# MT-HCCAR: Multi-Task Deep Learning with Hierarchical Classification and Attention-based Regression for Cloud Property Retrieval

### v1

@@ Abstract
In the realm of Earth science, effective cloud property retrieval, encompassing cloud masking, cloud phase classification, and cloud optical thickness (COT) prediction, remains pivotal. Traditional methodologies necessitate distinct models for each sensor instrument due to their unique spectral characteristics. Recent strides in Earth Science research have embraced machine learning and deep learning techniques to extract features from satellite datasets’ spectral observations. However, prevailing approaches lack novel architectures accounting for hierarchical relationships among retrieval tasks. Moreover, considering the spectral diversity among existing sensors, the development of models with robust generalization capabilities over different sensor datasets is imperative. Surprisingly, there is a dearth of methodologies addressing the selection of an optimal model for diverse datasets. In response, this paper introduces MT-HCCAR, an end-to-end deep learning model employing multi-task learning to simultaneously tackle cloud masking, cloud phase retrieval (classification tasks), and COT prediction (a regression task). The MT-HCCAR integrates a hierarchical classification network (HC) and a classification-assisted attention-based regression network (CAR), enhancing precision and robustness in cloud labeling and COT prediction. Additionally, a comprehensive model selection method rooted in K-fold cross-validation, one standard error rule, and two introduced performance scores is proposed to select the optimal model over three simulated satellite datasets OCI, VIIRS, and ABI. The experiments comparing MT-HCCAR with baseline methods, the ablation studies, and the model selection affirm the superiority and the generalization capabilities of MT-HCCAR.
Index Terms:
cloud property retrieval, remote sensing, deep learning, multi-task learning
I Introduction
The retrieval of cloud properties, denoting the estimation of diverse characteristics of clouds through the analysis of data acquired from remote sensing instruments, plays an essential role in atmospheric and environmental studies. This significance arises from the inherent role of clouds as integral components of the Earth’s atmospheric system, wielding substantial influence over the planet’s energy dynamics, climate regulation, and hydrological cycle[1]. The modulation of surface temperatures and the facilitation of precipitation by clouds underscore their paramount importance in sustaining ecosystems and shaping broader environmental processes.
Satellites serve as invaluable tools for global monitoring of cloud properties, offering an extensive data record of Earth’s imagery collected by various instruments since the latter part of the previous century. A prominent category of these instruments, commonly referred to as “imagers”, passively collect measurements of the Earth’s surface across multiple wavelength ranges, encompassing the ultraviolet (UV), visible (VIS), near-infrared (NIR), shortwave infrared (SWIR), and thermal infrared (TIR). The data obtained, primarily consisting of observations of reflected solar and/or emitted thermal radiation, undergo routine processing by algorithms. This processing transforms these measurements into geophysical quantities of interest, about both atmospheric and surface characteristics, in a process commonly termed “retrieval”. Routine retrieval of geophysical quantities is key to enhancing the comprehension of Earth’s climate [2]. At the end of such cloud retrieval process, each observed pixel is equipped with estimated values for many cloud properties including cloud mask (whether a pixel is clear or cloudy), cloud thermodynamic phase (liquid droplets or ice crystals), and cloud optical thickness (COT).
@@ I Introduction
One practical motivation for our work is NASA’s Plankton, Aerosol, Cloud, ocean Ecosystem (PACE) mission [3], which will launch in 2024 and advance science in its eponymous disciplines. Heritage cloud masking algorithms are not directly applicable to PACE’s main sensor, the Ocean Color Instrument (OCI) [4], due to different spectral bands compared to algorithms developed for existing sensors. OCI has similarities with some previous spaceborne imagers used for retrieval of cloud (and other) properties such as Moderate Resolution Imaging Spectroradiometer (MODIS) [5], Visible Infrared Imaging Radiometer Suite (VIIRS) [6], and Advanced Baseline Imager (ABI) [7] sensor types. However, two key differences are that 1) OCI has continuous hyperspectral coverage in the UV-NIR, plus discrete SWIR bands, while the others are only multi-spectral (up to a dozen or so discrete relevant bands); and 2) OCI lacks TIR bands. Some of the most commonly used cloud masking tests for those sensors are therefore not applicable, and adapting a subset of those tests would miss out on OCI’s unique abilities, so a new approach is warranted.
Numerous investigations have delved into the application of machine learning techniques for the extraction of cloud properties from satellite sensor observations. These inquiries encompass both singular cloud property retrievals [8, 9, 10, 11], as well as the concurrent retrieval of multiple cloud properties[12, 13, 14, 15]. Nevertheless, challenges and unresolved inquiries persist within the current landscape of machine learning-based cloud retrieval. First, the role of atmospheric domain knowledge, specifically the utilization of physical relationships between cloud mask and cloud phase, in conjunction with advanced machine learning techniques, remains unclear in enhancing the accuracy of cloud retrieval. Second, despite the deployment or design of numerous satellite sensors catering to analogous cloud retrieval tasks (expounded upon in Section III), the feasibility of employing a unified machine learning architecture to facilitate retrieval across different sensors and the generalization ability of such a machine learning model remain uncertain. Third, while there are various metrics to measure machine learning model performance, it remains ambiguous how an overarching ranking system can be established to select an optimal model, especially about determining whether the heightened complexity inherent in newly proposed models is justified in terms of their capabilities.
To address the above challenges, we propose an end-to-end Multi-Task learning architecture with Hierarchical Classification and Attention-assisted Regression (MT-HCCAR) for cloud masking, cloud thermodynamic phase (liquid or ice) prediction, and COT regression. The model consists of a Hierarchical Classification network for cloud masking and cloud phase prediction, and a Classification-assisted Attention-based Regression network (MT-HCCAR) for COT prediction. The classification task is a hierarchical classification (HC) sub-network to capture the hierarchical relationship between ice-crystal cloud and liquid cloud. The COT prediction task is a sub-network integrated with coarse cloud thickness classification (C) using attention mechanism (A) to further enhance regression (R) task accuracy. Our implementation is open-sourced at[16]. Our work’s major contributions are as follows.
  • We propose an end-to-end multi-task learning model MT-HCCAR for cloud masking, cloud phase classification, and COT regression. The model utilizes a hierarchical classification approach for cloud masking and cloud phase prediction. Meanwhile, to enhance COT regression accuracy, we use a cross-attention mechanism to learn similarities across the classification task and the regression task. Our extensive experiment results show performance improvement in all tasks using our approach than baseline models.
  • We conduct experiments on three satellite sensors (OCI, VIIRS, and ABI, introduced later) to compare different model performances across different satellite products. The results show the generalizability of our proposed model and we analyze our findings from different sensors’ performance.
  • To help identify the best model for practical usage, we design a novel model selection method that ranks the performance of different models across multiple metrics and multiple datasets.
The subsequent sections of this paper are structured as follows: Section II presents an overview of related work in the domains of cloud property retrieval and Multi-Task Learning (MTL) methods for remote sensing. Section III provides an in-depth exploration of the application background, the methodology employed for data simulation, and a precise articulation of our defined problem. Our proposed method, MT-HCCAR, is introduced in Section IV. The assessment metrics and outcomes of our experiments are elucidated in Section V. Section VI outlines our strategies for optimal model selection and presents the corresponding results. Lastly, Section VII encapsulates the paper with concluding remarks.
II Related work
Our study focuses on the application of multi-task learning on pixel-wise spectral features from satellite images. This section discusses the existing research on machine learning based cloud property retrieval methods and related work of multi-task learning applied to remote sensing data.
II-A Machine learning based cloud property retrieval
In the field of remote sensing, cloud properties encompass a diverse set of characteristics and parameters associated with clouds, which can be observed and measured from a distance using satellite or other remote sensing technologies. These properties include, among others, cloud mask, cloud (thermodynamic) phase, cloud vertical distribution, and COT. Machine learning approaches have demonstrated their efficacy in processing high-dimensional remote sensing data, automatically learning intrinsic features from spectrum data.
In recent years, various machine learning techniques, such as Random Forest (RF), Multi-Layer Perceptron (MLP), and Convolutional Neural Network (CNN), have been employed for retrieving different cloud properties. For instance, Liu et al. [8] utilized Artificial Neural Networks (ANN) and RF for cloud detection using spectral observations and reference labels from the Cloud-Aerosol Lidar with Orthogonal Polarization (CALIOP). Wang et al. [9] proposed an RF-based algorithm using passive spectral observations for cloud detection and phase classification across multiple instruments, evaluating their models against CALIOP and existing VIIRS cloud products. Giuffrida et al. [10] introduced CloudScout, a custom CNN deployed on the Hyperscout-2 nanosatellite payload for onboard cloud detection on hyperspectral images. He et al. [11] proposed a lightweight network, DABNet, for high-accuracy cloud detection based on CNN. Mohajerani et al.[17] presented Cloud-Net, a deep learning-based model using convolutional layers for cloud detection in Landsat 8 imagery. Shao et al. [18] proposed a multi-scale features CNN method for simultaneous classification of imagery into thick, thin, and non-cloud regions, outperforming other cloud detection methods. Jeppesen et al. [19] introduced RS-Net, a deep learning-based cloud detection algorithm for optical satellite imagery, achieving high performance even with a single spectral band. Segal-Rozenhaimer et al. [20] proposed a domain adversarial neural network (DANN) to enhance the transferability of cloud masks. Huang et al.[14] presented DAMA-WL, an end-to-end deep domain adaptation approach for cloud phase detection, incorporating domain alignment of heterogeneous active and passive satellite remote sensing data. Guo et al.[21] introduced CDnetV2, a CNN-based cloud detection method for remote sensing imagery with cloud-snow coexistence, utilizing attention-based modules for adaptive feature fusion. López-Puigdollers et al.[22] conducted a benchmark of deep learning models for cloud detection in multispectral images, including both rule-based approaches and machine learning approaches.
@@ II-A Machine learning based cloud property retrieval
While these machine learning approaches for cloud property retrieval predominantly leverage spectral features, the main limitations lie in two aspects: 1) many studies do not consider background knowledge, such as the task order (e.g., cloud mask prediction, cloud phase prediction, and COT prediction), and 2) several studies conduct different tasks independently, lacking parameter sharing between classification and regression tasks. These limitations underscore the need for more integrated and informed methodologies in cloud property retrieval studies.
II-B Multi-task learning for remote sensing data
Multi-task (MT) learning has emerged as a valuable approach in Earth science and remote sensing applications, facilitating the retrieval of related products from shared features. Recent studies have demonstrated the efficacy of the multi-task learning paradigm in jointly enhancing the performance of diverse tasks.
Ilteralp et al. [23] proposed an MT Convolutional Neural Network (CNN) architecture capable of leveraging unlabeled satellite imagery to predict Chl-a values, incorporating an auxiliary classification task. The proposed MT learning model was compared against alternative approaches, including its single-task regression version, support vector regression, Multi-Layer Perceptron (MLP), morphology-based spectral-spatial approach (EFAL), CNN, and statistical spectral band characteristics known as the band ratio. Zhang et al. [24] introduced MTDR-Net, an MT-driven and reconfigurable network designed to improve cloud detection in regions with coexisting cloud and snow cover in very high-resolution remote sensing images. This model employs an MT gradient flow guidance module (MTGFG) to establish a feature-sharing strategy across different tasks. Chen et al. [25] proposed an MT learning method based on Long Short-Term Memory (LSTM) with an attention mechanism to classify imbalances in wind turbine blades caused by ice accretion. Sarafian et al. [26] presented a meteorology-based model for dust event forecasting, utilizing an MT framework for predicting regional PM10 fields and local in situ measured PM10 values. Kim et al. [27] proposed MT CNN and ConvLSTM models for simultaneous forecasting of pixel-wise sea ice concentration, considering both spatial sea ice concentration (SIC) images and temporal sea ice extent (SIE) values. Qiu et al. [28] developed an end-to-end CNN with a shared backbone network for shared feature learning, attention modules for task-specific feature learning for human settlement extent (HSE) regression, and local climate zone (LCZ) classification. Hanna et al. [29] proposed a U-Net-based architecture to simultaneously predict the pixel area covered by plumes, the type of fired fuel, and the power generation rate from a single-satellite image of a power plant. Ahn et al. [30] presented a systematic approach to forecasting regional wave heights based on different LSTM structures. Zhao et al. [31] introduced a CNN-based framework that combines self-supervised learning and scene classification to improve feature extraction and generalization abilities.
@@ II-B Multi-task learning for remote sensing data
There are also studies conducting ML learning specifically for cloud property retrieval. Yang et al.[12] developed an MLP-based method to retrieve cloud macrophysical parameters (cloud mask, cloud top temperature, and cloud top height) using Himawari-8 satellite data. Wang et al.[13] proposed TIR-CNN based on the U-Net model to retrieve cloud properties including cloud mask, COT, effective particle radius (CER), and cloud top height (CTH) from thermal infrared radiometry. The architecture consists of encoding and decoding layers, convolutional blocks, batch normalization layers, and leaky rectified linear units. The results of applying the model to thermal infrared spectral data from MODIS are used to compare model performance for daytime and nighttime data.
@@
While these above studies offer valuable insights that are beneficial to our design, our work cannot be compared with most of them directly via experiments. It is because of the distinctive nature of our model’s output, which manifests at the pixel level rather than the image level. Consequently, architectures employing convolutional layers, such as U-net architecture, are ill-suited for our study’s requirements.
III Problem statement and data simulation
In this section, we first present the contextual background of the application which includes a forward simulation process and an inverse retrieval process. Then, we delve into a comprehensive exposition of the employed data simulation method in Section III-A, and define our cloud property retrieval problem based on the generated simulation data in Section III-B.
The forward simulation and inverse retrieval processes are characterized as two opposite operations within the context of cloud property analysis. In the forward simulation, spectral bands and physical variables are generated as outputs using cloud properties as input parameters. Conversely, cloud property retrieval involves utilizing spectral features obtained from satellite observations as input, with the resultant output comprising cloud properties such as cloud mask, cloud phase, and COT values. To provide a more explicit formulation, the forward simulation is represented as the mathematical process:
$\mathbf{X}=f(\mathbf{L}_{cm},\mathbf{L}_{cph},\mathbf{Y}_{COT},...)$ (1)
@@ III Problem statement and data simulation
, where $\mathbf{X}=[\mathbf{x_{1}},\mathbf{x_{2}},\ldots,\mathbf{x_{M}}]$ denotes the spectral and physical features, $f(\cdot)$ represents the simulation algorithm, and $\mathbf{L}_{cm},\mathbf{L}_{cph},\mathbf{Y}_{COT}$ are the labels corresponding to cloud mask, cloud phase, and COT values, respectively. Additional properties are also used as inputs (see Section III-A), but they are not the focus of this work.
Conversely, the inverse retrieval process can be expressed as:
$\mathbf{\hat{L}}_{cm},\mathbf{\hat{L}}_{cph},\mathbf{\hat{Y}}_{COT}=g(\mathbf{X})$
, where $[\mathbf{\hat{L}}_{cm},\mathbf{\hat{L}}_{cph},\mathbf{\hat{Y}}_{COT}]$ represents the predicted values of cloud mask, cloud phase, and COT, as determined by the retrieval algorithm $g(\cdot)$ given the input spectral and physical features $\mathbf{X}$. Figure 1 illustrates how these two processes operate in our study and the colors of the three sensors show their wavelength band differences.
Figure 1: Illustration of forward simulation and inverse retrieval. Forward simulation uses cloud properties as input parameters to generate spectral bands and physical variables as outputs. Conversely, inverse retrieval entails using spectral features obtained from satellite observations as inputs to generate cloud properties such as cloud mask, cloud phase, and COT values.
III-A Forward radiative transfer simulation
Given that OCI has not been launched, simulated datasets for OCI, VIIRS, and ABI are employed for assessing sensors and architectures against a standardized source. Radiative transfer (RT) models play a crucial role by capturing optical processes like scattering and absorption. These models map geophysical parameters, representing surface and atmospheric properties, to spectral intensities at specific detection points—essentially serving as the intermediary link between geophysical quantities and observations made by space-borne sensors.
We cover from ultraviolet (UV) to shortwave infrared (SWIR) and then convolve these hyperspectral simulations with the solar spectrum from [32] and sensor relative spectral response functions for OCI (nominal pre-launch) [4], ABI [7], and VIIRS [6] in order to generate the spectral top of atmosphere reflectance signal that the instruments would observe. This provides our simulated observations and reference truth (i.e., cloud classification and COT), with band centers simulated for each sensor as below:
  • OCI: 233 bands in total; 226 from the hyperspectral detectors from 350-890 nm (with 2.5 nm spacing between bands) plus seven discrete NIR/SWIR bands centered near 940, 1040, 1250, 1378, 1620, 2130, and 2260 nm.
  • VIIRS: 10 bands in total, centered near 412, 445, 488, 555, 672, 865, 1240, 1380, 1610, and 2250 nm.
  • ABI: 6 bands in total, centered near 471, 640, 860, 1370, 1600, and 2200 nm.
Each ABI band has a close analog in VIIRS, and likewise, each VIIRS band has a close analog in OCI, so the trio represent a set of increasing complexity. Both ABI and VIIRS also have thermal infrared (TIR) bands, but these are excluded as the primary motivation for this work is the development of a new methodology for OCI (which does not), and the application to ABI and VIIRS demonstrates the broader applicability of our new architectures.
Inputs for the RT simulation include cloud (phase, microphysical and optical properties, and vertical structure), atmospheric profile (aerosol, gas, Rayleigh scattering, temperature, and pressure), surface (spectral reflectance), and observation (solar and view angles) properties. An ensemble of 250,000 simulations, parameterized by independent samples from realistic distributions of these geophysical properties, form the labeled dataset for model training and evaluation. Distributions and further specifics of the simulations are described in Sayer et al. [33], with two deviations: 1) that study consisted of purely cloud-covered parameterizations; here, $1/3$ of simulations corresponded to cloud-free conditions, and the rest to single-layer clouds, and 2) that study considered a narrow spectral range so assumed a spectrally-flat surface albedo; here, we use 20-class extended International Geosphere-Biosphere Program (IGBP) albedo system [34]. These 20 classes correspond to 16 types of land surface spectra (1 desert, 15 non-desert), 1 water surface, and 3 snow/sea ice surfaces. We sampled such that a given simulation had a 1/34 chance of being drawn from each of the 16 “land” spectra, a 9/34 chance of using the “water” spectrum, and a 3/34 of being drawn from each of the “snow/ice” spectra, and for network training classes collapse the 20 spectra into these four main categories (water; snow/ice; desert; non-desert land). The simulation does not include multi-layer or mixed-phase cloud systems, as most satellite retrieval processing algorithms assume a single-layer single-phase cloud system. However, future work could extend the network to be trained (or just evaluated) on an ensemble including such systems.
We emphasize that the software simulates individual point observations (i.e., a pixel) which means that (as is standard for most cloud retrievals in Earth science) the network operates on single pixels at once as opposed to larger scenes (i.e., no spatial context). It would be computationally prohibitive to simulate a globally representative ensemble of satellite scenes, particularly the inclusion of 3D radiative transfer effects. As a result, existing works involving 3D effects[35, 36] have typically been limited to a few case studies.
III-B Problem definition for inverse cloud property retrieval

### v2

@@ Abstract
In Earth science, accurate retrieval of cloud properties including cloud masking, cloud phase classification, and cloud optical thickness (COT) prediction is essential in atmospheric and environmental studies. Conventional methods rely on distinct models for each sensor due to their unique spectral characteristics. Recently, machine/deep learning has been embraced to extract features from satellite datasets, yet existing approaches lack architectures capturing hierarchical relationships among tasks. Additionally, given the spectral diversity among sensors, developing models with robust generalization capabilities remains challenging for related research. There is also a notable absence of methods evaluated across different satellite sensors. In response, we propose MT-HCCAR, an end-to-end deep learning model employing multi-task learning. MT-HCCAR simultaneously handles cloud masking, cloud phase retrieval (classification tasks), and COT prediction (a regression task). It integrates a hierarchical classification network (HC) and a classification-assisted attention-based regression network (CAR), enhancing precision and robustness in cloud labeling and COT prediction. Experimental evaluations, including comparisons with baseline methods and ablation studies, demonstrate that MT-HCCAR achieves optimal performance across various evaluation metrics and satellite datasets.
Abstract
This additional material offers supplementary details of experiments discussed in the main paper published in ECML PKDD 2024.
  • Section A: Experiment Settings of Ablation Study (additional discussion to support the beginning of Section 5 and Table 1 in the main paper).
  • Section B: Cross-Validation and Model Selection (additional discussion supporting Section 5.4 in the main paper).
Keywords:
Cloud property retrieval Remote sensing Deep learning.
1 Introduction
Clouds are integral components of the Earth system, wielding substantial influence over the planet’s energy dynamics, climate regulation, and the hydrological cycle [24]. Satellites have long been an indispensable tool to help us understand our environment. A prominent category of these instruments, commonly referred to as imagers, passively collect measurements of the Earth across various combinations of ultraviolet (UV), visible (VIS), near-infrared (NIR), shortwave infrared (SWIR), and thermal infrared (TIR) wavelength ranges. These measurements of reflected solar and/or emitted thermal radiation in different spectral bands undergo routine processing by algorithms to convert them into geophysical parameters of interest (atmospheric and surface characteristics) in a process known as “retrieval”. This is typically done pixel-by-pixel for each satellite image. For clouds, several key properties are targeted, including a cloud mask (which distinguishes cloud-covered pixels from cloud-free ones), thermodynamic phase (indicating whether the cloud comprises liquid water droplets or ice crystals), and cloud optical thickness (COT), which is a measure of both the amount of light scattered by a cloud and the quantity of liquid or ice within it. The routine retrieval of geophysical parameters is essential for advancing our understanding of Earth’s climate [13].
@@ 1 Introduction
One practical motivation for our work is NASA’s Plankton, Aerosol, Cloud, ocean Ecosystem (PACE) mission [28], which launched in February 2024. Existing NASA cloud masking algorithms are not directly applicable to PACE’s main sensor, called OCI [5], due to different spectral bands compared to algorithms developed for existing sensors. OCI has similarities with some previous spaceborne imagers such as MODIS [3], VIIRS [6], and ABI [1] sensor types. However, two key differences are that 1) OCI has continuous hyperspectral coverage in the UV-NIR, plus discrete SWIR bands, while the others are only multi-spectral (up to a dozen or so discrete relevant bands); and 2) OCI lacks TIR bands. Some of the most commonly used cloud masking tests for those sensors are therefore not applicable, and adapting a subset of those tests would miss out on OCI’s unique abilities, so a new approach is warranted.
Numerous studies have explored machine learning techniques to extract cloud properties from satellite sensor data. These include both retrieval of a single cloud property [17, 25, 12], and simultaneous retrievals of multiple cloud properties [26, 14]. However, challenges persist in this area. To begin with, it’s unclear how incorporating atmospheric domain knowledge, such as physical relationships between cloud properties, alongside advanced machine learning techniques, enhances retrieval accuracy. Furthermore, despite the deployment of numerous satellite sensors for similar cloud retrieval tasks (as discussed in Section 3), the generalizability, especially from an Earth science perspective, of employing a unified machine learning architecture across different sensors remains uncertain.
To address the above challenges we introduce MT-HCCAR, an end-to-end Multi-Task Learning (MTL) model for cloud masking, cloud phase prediction, and COT regression. Our contributions are fourfold. First, we incorporate hierarchical classification to capture the hierarchical relationship between cloud masking and cloud phase classification, enhancing prediction performance as demonstrated in the comparison experiment with baseline methods and ablation study. Second, we employ a cross-attention module to improve COT regression accuracy by leveraging similarities across the classification and regression networks. Third, we conduct quantitative experiments to demonstrate the advantages of our MT-HCCAR model over state-of-the-art baselines and its generalizability across three different satellite sensors. Finally, we perform quantitative and qualitative evaluation of performance variations across sensors from an Earth science perspective. Our implementation is publicly available at GitHub [2].
2 Related Work
@@
In recent years, various machine learning techniques, such as Random Forest (RF), Multi-Layer Perceptron (MLP), and Convolutional Neural Network (CNN), have been employed for retrieving different cloud properties. Among them, most of the work targets cloud detection [17, 25], cloud phase [11], and cloud thickness [21]. While these machine learning approaches for cloud property retrieval predominantly leverage spectral features, the main limitations lie in two aspects: 1) many studies do not consider background knowledge, such as the task order (e.g., cloud mask prediction, cloud phase prediction, and COT prediction), and 2) several studies conduct different tasks independently, lacking knowledge sharing between classification and regression tasks. These limitations underscore the need for more integrated and informed methodologies in cloud property retrieval studies.
@@
Multi-task learning (MTL) is proving valuable in Earth science and remote sensing by jointly enhancing performance across diverse remote-sensing tasks through shared features. The studies have shown that classification and regression tasks can be implemented together to improve model performance [15, 7]. There are also studies conducting ML learning specifically for cloud property retrieval. Yang et al. [29] developed an MLP-based method to retrieve cloud macrophysical parameters (cloud mask, cloud top temperature, and cloud top height) using Himawari-8 satellite data. Wang et al. [26] proposed TIR-CNN based on the U-Net model to retrieve cloud properties including cloud mask, COT, effective particle radius (CER), and cloud top height (CTH) from thermal infrared radiometry. The architecture consists of encoding and decoding layers, convolutional blocks, batch normalization layers, and leaky Rectified Linear Units (ReLU). The results of applying the model to thermal infrared spectral data from MODIS are used to compare model performance for daytime and nighttime data.
@@
While these above studies offer valuable insights that are beneficial to our design, our work cannot be compared with most of them directly via experiments. This is because the OCI data we use does not have the same spectral bands, such as TIR bands, with those methods used. Consequently, we will not compare our model with architectures employing convolutional layers, such as CNN.
@@
Figure 1: Illustration of forward simulation and retrieval. $\mathbf{x_{M}}$: the spectral and physical features. $f(\cdot)$, $g(\cdot)$: the RT simulation algorithm and the retrieval algorithm. $\mathbf{L}_{cm},\mathbf{L}_{cph},\mathbf{Y}_{COT}$: the labels of cloud mask, cloud phase, and COT values, respectively.
3 Problem Statement and Data Simulation
Cloud property determination requires two operations: simulation and retrieval. Simulation is a forward process mapping from the Earth’s state (cloud and surface properties) to satellite direct observations (the reflected light signal the satellites see). Retrieval is an inverse process, determining the geophysical quantities from the satellite observations. Figure 1 illustrates these two processes with colors showing differences in the wavelength band of the three sensors.
3.1 Radiative Transfer Simulation
The simulation process provides training data for our cloud property retrieval model. This study was conducted before the launch of PACE, and the spacecraft has not entered routine operations. Therefore, simulated datasets for OCI, VIIRS, and ABI were utilized to evaluate the sensors and architectures against a standardized source, enabling direct comparisons. This approach ensures consistency as the satellites are tested on identical simulated scenes with the same viewing geometries, which would not be feasible using real data.
Radiative transfer (RT) models describe the scattering and absorption processes that affect the propagation of light through Earth’s atmosphere and surface. RT is the forward model to map from surface and atmospheric properties to spectral intensities seen by the space-borne sensors. We account for realistic variations in the properties of clouds (phase, microphysical, optical, and vertical structure), the atmospheric profile (aerosol, gas, Rayleigh scattering, temperature, and pressure), surface reflectance, and solar/observation geometry. Further details are as in Sayer et al. [20], with two modifications: 1) simulating cloud-free conditions as well as single-layer clouds, and 2) utilizing 20 different surface reflectance classes included in the libRadtran RT model [10] (including various land surface spectra, water, and snow/sea ice). The simulation does not include multi-layer or mixed-phase cloud systems, aligning with most satellite retrieval processing algorithms. However, future work may extend the network to include such systems. We generate a dataset of 250,000 simulations for model training and evaluation. We simulate from the UV to SWIR and convolve these simulations with the solar spectrum from [9] and sensor relative spectral response functions for OCI [5], ABI [1], and VIIRS [6] in order to generate the spectral top of atmosphere reflectance signal that the instruments would observe. This provides our simulated observations and reference truth (cloud classification and COT), along with band centers for each sensor. For OCI, there are 233 bands in total, including 226 hyperspectral bands ranging from 350 to 890 nm with 2.5 nm spacing, and seven discrete NIR/SWIR bands centered near 940, 1040, 1250, 1378, 1620, 2130, and 2260 nm. VIIRS has 10 bands centered near 412, 445, 488, 555, 672, 865, 1240, 1380, 1610, and 2250 nm. ABI has 6 bands centered near 471, 640, 860, 1370, 1600, and 2200 nm, respectively.
Each ABI band has a close analog in VIIRS, and similarly, each VIIRS band has a counterpart in OCI, thus forming a trio representing increasing complexity. Another commonly-used sensor, MODIS, has similar spectral coverage to VIIRS so we do not include it. Both ABI and VIIRS have TIR bands, but these are omitted from consideration as our primary focus is on developing a new methodology for OCI, which lacks TIR bands. Application to ABI and VIIRS demonstrates the broader applicability of our new architectures.
3.2 Cloud Property Retrieval